# Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model

### v2

@@ Abstract
The growth of quantum uncertainty at different stages of quantum scrambling can be characterized by out-of-time-order correlation functions (OTOCs). For instance, quantum uncertainty around the Ehrenfest time grows exponentially [1, 2, 3] in the semiclassical limit for quantum chaotic systems. The calculation of OTOCs for quantum chaotic systems is quite challenging with explicit results known only for certain quantum maps [3], a kicked rotor [2], and a particle in a disordered potential [1]. For integrable systems the growth is typically power-law. Reported exponential growth [4, 5, 6, 7, 8, 9] in certain integrable systems requires [10, 11] choosing initial conditions around unstable fixed points so that the evolution of wave functions mimics that of quantum chaotic systems for short times due to the exponential separation of classical trajectories.
For many-body quantum chaotic systems, the Lyapunov exponent was first computed analytically by Kitaev [12] in the low-temperature limit of the SYK model [12, 13, 14, 15, 16, 17, 18] consisting of $N$ fermions with $q$-body random interactions of infinite range. For $q>2$, the dynamics is quantum chaotic at all timescales [12, 19, 18] with a Lyapunov exponent that saturates a universal bound on chaos [20]. Other aspects of the OTOC in the Hermitian SYK model have been studied analytically in Refs. [18, 21, 22, 23, 24, 25] and numerically, for $q>2$, in Refs. [26, 27] where it was necessary to reach $N\geq 50$ Majoranas in order to reproduce the mentioned saturation of the Lyapunov exponent. The central role of the OTOC in the description of quantum scrambling dynamics has triggered a flurry of activity [11] in different fields. For instance, the OTOC has been studied for a random matrix Hamiltonian [28, 29, 30], for many-body bosonic systems  [31, 32], random circuits [33, 34, 35] and Jackiw-Teitelboim gravity [36, 37, 38, 39, 40].
@@
Model and analytic calculation of the OTOC.—We consider an integrable Hermitian SYK Hamiltonian coupled to a Markovian bath at inverse temperature $\beta$. In this case, the dynamics is described by the Lindblad formalism [59, 57] and depends on the choice of jump operators and temperature-dependent couplings. We start our analysis with the infinite-temperature case, $\beta=0$, where calculations are especially simple. The evolution of the density matrix is governed by the Lindblad equation,
@@
where $H=i\sum_{i<j}J_{ij}\chi_{i}\chi_{j}$ is the $q=2$ SYK Hamiltonian with $\{\chi_{i},\chi_{j}\}=\delta_{ij}$, $L_{i}=\sqrt{\mu}\chi_{i}$, and $J_{ij}=-J_{ij}$ are Gaussian random numbers of zero mean and variance $\langle J_{ij}^{2}\rangle=J^{2}/N$. We set $J=1$, so all results are in units of $J$. The steady state is the identity, namely, a thermofield double state at infinity temperature. We aim to probe the dynamics by the study of the growth in time of quantum uncertainty represented by the square of anti-commutators, $C(t)=\frac{2^{-N/2}}{N(N-1)}\sum_{i\neq j}\langle\{\chi_{i}(t),\chi_{j}\}^{2}\rangle\equiv 2F(t)+2I(t)$, where $\chi_{k}$ without explicit arguments stands for its value at $t=0$, $F(t)=2^{-N/2}(N(N-1))^{-1}\sum_{m\neq n}\langle\chi_{m}(t)\chi_{n}\chi_{m}(t)\chi_{n}\rangle$ and $I(t)=C(t)/2-F(t)$. $F(t)$ is the OTOC, which for quantum chaotic systems captures the exponential growth of the quantum uncertainty around the Ehrenfest time. The bra-kets denote a trace and an average over the random SYK couplings. We focus on $F(t)$ since the calculation for $I(t)$ is similar. To determine the $\mu$-dependence of $F(t)$, we can restrict ourselves to two specific Majoranas, $\chi_{1}$ and $\chi_{2}$, and employ the quantum regression theorem for OTOCs [62]:
$\frac{dF}{dt}=$ $-2i\langle[H,\chi_{1}(t)]\chi_{2}\chi_{1}(t)\chi_{2}\rangle-\mu N\langle\chi_{1}(t)\chi_{2}\chi_{1}(t)\chi_{2}\rangle$
@@
Note the extra minus sign in the last term [compared to Eq. (1)], which arises in the adjoint Lindblad evolution with fermionic jump operators [63]. Using $\sum_{n}\chi_{n}\chi_{k}\chi_{n}=-(N/2-1)\chi_{k},$ it is straightforward to show that the $\mu$ dependence of the OTOC factorizes, $F(t,\mu)=e^{-2\mu t}F(t,\mu=0)$ so the dependence on $C(t)$ factorizes as well.
We now turn to the finite temperature case. The proof of factorization of the $\mu$-dependence at finite temperature, presented in Appendix A, follows along the same lines as for $\beta=0$. A finite-temperature steady state $\rho=\exp(-\beta H)/Z$, with $Z={\rm Tr}\exp(-\beta H)$ can be prepared by a judicious choice of jump operators with temperature-dependent couplings [64] (see Appendix A). In this case, we define the quantum uncertainty as
$C(t,\beta,\mu)$ $=\sum_{i\neq j}\frac{\langle\{\chi_{i}(t)\rho^{1/4},\chi_{j}\rho^{1/4}\}^{2}\rangle}{N(N-1)}$
$=e^{-2\mu t}(F(t,\beta)+I(t,\beta)),$ (3)
where $F$ and $I$ are defined with the corresponding $\rho^{1/4}$ insertions  [18] with respect to the $\beta=0$ analogues. As a consequence, it is only necessary to compute analytically $C(t)$ for a Hermitian $q=2$ SYK model at finite temperature.
In order to proceed, we choose a basis in which $H$ is diagonal, $H=i\sum_{k=1}^{N/2}\lambda_{k}\widetilde{\chi}_{2k-1}\widetilde{\chi}_{2k}$, where $\widetilde{\chi}_{i}=S_{ik}\chi_{k}$, with $S_{ik}$ an orthogonal matrix, and $\lambda_{k}$ are the eigenvalues of the antisymmetric real couplings $J_{ij}$. In this new basis, the OTOC is given by
@@
[…]
It is now necessary to carry out the averages over the orthogonal matrices $S_{ij}$. This is possible by employing the relation [65] $\langle S_{1k}S_{2m}S_{1p}S_{2q}\rangle=(-\delta_{km}\delta_{qp}-\delta_{kq}\delta_{mp}+(N+1)\delta_{kp}\delta_{mq})/\mathcal{N}$, where ${\cal N}=N(N-1)(N+2)$, which results in three different contributions to $F(t)$. The first two are equal because of the reflection symmetry of the spectrum. The OTOC can be further simplified using $\widetilde{\chi}_{k}^{2}=1/2$ after commuting the four Majorana operators through the evolution operator and the density matrix. The resulting sums are expressed in terms of $\lambda_{k}$ by using the properties of the trace over the many-body states. The final finite-$N$ result for $C(t,\beta)$ Eq. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model) is
@@
Analytic expressions to orders $1/N$ and $1/N^{2}$.—The above expressions simplify by keeping only the leading $1/N$ correction, replacing the sums with integrals and noticing that the spectral density of the eigenvalues $\lambda_{k}$ is $\rho(\lambda)=(1/\pi)\sqrt{1-\lambda^{2}/4}$ [66]. For simplicity, we henceforth restrict ourselves to $\beta=0$. Considering only the leading $1/N$ correction, $C(t,0)\equiv C(t)$ Eq. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model), with $F(t,0)$ and $I(t,0)$ given by Eqs. (5), (6), simplifies to
@@
where $J_{1}(t)$ is a Bessel function of first kind and time is measured in units of $J=1$. The upper plot in Fig. 1 demonstrates the convergence of the numerical results of $C(t)$ obtained from Eqs. (5) and (6) to the analytic prediction Eq. (8). Analogous analytic results for finite $\beta$ are rather cumbersome. In Appendix A, we have checked that finite temperature effects do not change qualitatively the dynamics, at least for sufficiently long times. A consequence of Eq. (8) is that, for $t\ll 1$, $C(t)$ grows quadratically as $t^{2}$. For $t\gg 1$, the system approaches a steady state in a power-law fashion with superimposed oscillations. This is in contrast with quantum chaotic systems for which [3, 11] this approach is exponential with no oscillations resulting in full equilibration. The effect of the bath is an overall exponential decay of $C(t)$ that, as expected, suppresses quantum scrambling. For a sufficiently weak coupling, $\mu\ll 1$, the system approaches the steady state in two stages: first, the mentioned power-law, and only for longer times $\sim 1/\mu$ the exponential suppression of $C(t)$.
Figure 1: Time evolution of the quantum uncertainty $C(t)$ with $t$ in units of $J=1$ for $\mu=\beta=0$. Top: The blue ($N=20$), red ($N=200$), and green ($N=2000$) curves correspond to $C(t)$ given in Eq. (5) and (6) and $C_{N}=(N+1)/[(N-1)(N+2)]$. They converge to the $1/N$ prediction Eq. (8), with $C_{N}=1/N$, only for sufficiently large $N$. In both cases, $C_{N}$ is given by the time independent part of $C(t)$. Bottom: Comparison between the analytical exact $C(t)$ Eqs. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model), (5), (6) in terms of the eigenvalues of $J_{ij}$ (purple circles) with the simple analytic $C(t)$ including up to $1/N^{2}$ corrections for $N=200$ on a linear scale. The black curve stands for the $1/N$ result, Eq. (8), while the green curve stands for the $1/N+1/N^{2}$ result, Eqs. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model) and (11). The agreement between of analytical exact and the $1/N+1/N^{2}$ results is excellent for all timescales. Deviations from the $1/N$ expression occur at $t\sim\sqrt{N}$. The Heisenberg time, signaling saturation, is at $t\sim 2N$.
@@
where we have reinstated $\cal N$ in order to account for additional $1/N^{2}$ effects, $\theta$ is the Heaviside step function, and $N^{2}K_{c}(t)=\langle|\sum_{k=1}^{N}\exp(it\lambda_{k})|^{2}\rangle-|\langle\sum_{k=1}^{N}\exp(it\lambda_{k})\rangle|^{2}$ is the connected spectral form factor (SFF).
@@
For times $t\gtrsim\sqrt{N}$, $K_{c}(t)\sim 2t/\pi N$ and then gradually saturates to $1/N$ as the Heisenberg time $t\sim 2N$ is approached. We stress that since the spectral density is not constant, the time dependence of Eq. (11) has clear deviations from linear behavior before the Heisenberg time, which translates into similar deviations in $C(t)$. Interestingly, in the range of times of interest, $2N>t>\sqrt{N}$, a simple inspection of Eq. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model) reveals that $C(t)$ is controlled by $K_{c}(t)$ despite it being of the order $1/N^{2}$ and therefore subleading in the $1/N$ expansion. The reason for that is that the time dependence of the leading $1/N$ terms approaches zero in a power-law fashion so that for $t>\sqrt{N}$ they become smaller than $K_{c}(t)$ which increases linearly in time. Crucially, the sign of the SFF term is negative, which leads to a local maximum in the OTOC not related to the small oscillations. Therefore, the OTOC decreases with time for $t>\sqrt{N}$, which corresponds to a decrease of quantum scrambling due to the discreteness of the spectrum captured by the SFF.
@@
The results depicted in the top panel of Fig. 1 fully confirm this picture. The analytical expression for $C(t)$ Eq. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model), which includes $1/N$ and $1/N^{2}$ corrections, agrees at all times scales with the exact $C(t)$ Eqs. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model), (5), (6), at $\beta=0$, in terms of the eigenvalues of $J_{ij}$, which must be computed numerically. Substantial deviations are observed for long times if only the leading $1/N$ correction Eq. (8) is taken into account. The effect of the bath is to suppress scrambling exponentially.
@@
We have thus found that, for $t\gtrsim\sqrt{N}$, $C(t)$ is controlled by the SFF of the random couplings $J_{ij}$, which has the expected linear behavior observed in quantum chaotic systems for times $t>\sqrt{N}$ but not too close to the Heisenberg time. This does not mean that the $q=2$ SYK is many-body quantum chaotic. While single-particle observables may show quantum chaos, the many-body dynamics are still integrable because the Majoranas are not interacting. For instance, the level statistics is Poisson. Moreover, the timescales involved are polynomial in $N$ which is typical of integrable systems. The results of Appendix A show that finite $\beta$ effects do not change qualitatively this picture.
@@
Figure 2: Top: $C(t)$ with $t$ in units of $J=1$, for $\mu=\beta=0$. $C(t)$ Eq. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model) is computed using exact diagonalization for $N=24$ (red) with $\sim 10^{6}$ disorder realizations. The exact diagonalization result shows almost no difference with respect to the exact analytic result (black) Eqs. (5), (6), and and also an excellent agreement with the $1/N+1/N^{2}$ simple expression (blue) Eq. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model). Substantial deviations are observed if only the leading $1/N$ correction (green) Eq. (8) is kept. Bottom: $C(t)$ for $\mu=0.15$, $N=18$ and $\beta=0$. We combine exact diagonalization with the quantum trajectory method [69, 70]. For each point in time, we use $200$ quantum trajectories, a time step of $dt=0.01$, and at least $5\times 10^{5}$ disorder realizations. The exponential approach to the steady state for $\mu\neq 0$ is correctly reproduced by the quantum trajectory method and we find good agreement with the $1/N$ semiclassical result.

### v3

@@ Abstract
I Introduction
The growth of quantum uncertainty at different stages of quantum scrambling can be characterized by out-of-time-order correlation functions (OTOCs). For instance, quantum uncertainty around the Ehrenfest time grows exponentially [1, 2, 3] in the semiclassical limit for quantum chaotic systems. The analytical calculation of OTOCs even for single-particle quantum chaotic systems is quite challenging with relatively few explicit calculations known [1, 3, 2]. For integrable systems, the growth is typically power-law. Reported exponential growth [4, 5, 6, 7, 8, 9] in certain integrable systems requires [10, 11] choosing initial conditions around unstable fixed points so that the state evolution mimics that of quantum chaotic systems for short times.
For many-body quantum chaotic systems, the Lyapunov exponent was first computed analytically by Kitaev [12] in the low-temperature limit of the SYK model [12, 13, 14, 15, 16, 17, 18] consisting of $N$ fermions with $q$-body random interactions of infinite range. For $q>2$, the dynamics is quantum chaotic at all timescales [12, 19, 18] with a Lyapunov exponent that saturates a universal bound on chaos [20]. Other aspects of the OTOC in the Hermitian SYK model have been studied analytically in Refs. [18, 21, 22, 23, 24, 25] and numerically, for $q>2$, in Refs. [26, 27] where it was necessary to reach $N\geq 50$ Majoranas in order to reproduce the mentioned saturation of the Lyapunov exponent. The central role of the OTOC in the description of quantum scrambling has triggered a flurry of activity [11] in different fields: the OTOC has been studied for a random matrix Hamiltonian [28, 29, 30], for many-body bosonic systems [31, 32], random circuits [33, 34, 35] and Jackiw-Teitelboim gravity [36, 37, 38, 39, 40].
@@ I Introduction
II Model and analytic calculation of the OTOC
We consider an integrable Hermitian SYK Hamiltonian coupled to a Markovian bath at inverse temperature $\beta$. In this case, the dynamics is described by the Lindblad formalism [59, 57] and depends on the choice of jump operators and temperature-dependent couplings. We start our analysis with the infinite-temperature case, $\beta=0$, where calculations are especially simple. The evolution of the density matrix is governed by the Lindblad equation,
@@ II Model and analytic calculation of the OTOC
where $H=i\sum_{i<j}J_{ij}\chi_{i}\chi_{j}$ is the $q=2$ SYK Hamiltonian with $\{\chi_{i},\chi_{j}\}=\delta_{ij}$, $L_{i}=\sqrt{\mu}\chi_{i}$, and $J_{ij}=-J_{ij}$ are Gaussian random numbers of zero mean and variance $\langle J_{ij}^{2}\rangle=J^{2}/N$. We set $J=1$, so all results are in units of $J$. The steady state is the identity, namely, a thermofield double state at infinity temperature. We aim to probe the dynamics by the study of the growth in time of quantum uncertainty represented by the square of anti-commutators,
$C(t)=\frac{1}{N(N-1)}\sum_{i\neq j}\langle\{\chi_{i}(t),\chi_{j}\}^{2}\rangle\equiv 2F(t)+2I(t),$ (2)
where $\chi_{k}$ without explicit arguments stands for its value at $t=0$, and
$F(t)$ $=$ $\frac{1}{N(N-1)}\sum_{m\neq n}\langle\chi_{m}(t)\chi_{n}\chi_{m}(t)\chi_{n}\rangle,$
$I(t)$ $=$ $C(t)/2-F(t).$ (3)
Here, $F(t)$ is the OTOC, which for quantum chaotic systems captures the exponential growth of the quantum uncertainty around the Ehrenfest time. The bra-kets denote an average over all states and an average over the random SYK couplings. We focus on $F(t)$ since the calculation for $I(t)$ is similar. To determine the $\mu$-dependence of $F(t)$, we can restrict ourselves to two specific Majoranas, $\chi_{1}$ and $\chi_{2}$, and for $q=2$ one can easily derive the evolution equation [62]:
$\frac{dF}{dt}=$ $2i\langle[H,\chi_{1}](t)\chi_{2}\chi_{1}(t)\chi_{2}\rangle-\mu N\langle\chi_{1}(t)\chi_{2}\chi_{1}(t)\chi_{2}\rangle$
@@
Note the extra minus sign in the last term [compared to Eq. (1)], which arises in the adjoint Lindblad evolution with fermionic jump operators [63],
$\frac{d\chi_{i}(t)}{dt}$ $=i[H,\chi_{i}(t)]$ (5)
$-\mu\sum_{n}\chi_{n}^{\dagger}\chi_{i}(t)\chi_{n}-\frac{\mu}{2}\sum_{n}\{\chi_{n}^{\dagger}\chi_{n},\chi_{i}(t)\}$
when deriving the evolution equation (II) with $\chi_{n}^{\dagger}=\chi_{n}$. Using
$\sum_{n}\chi_{n}\chi_{k}\chi_{n}=-(N/2-1)\chi_{k},$ (6)
it is straightforward to show that the $\mu$ dependence of the OTOC factorizes,
$F(t,\mu)=e^{-2\mu t}F(t,\mu=0),$ (7)
so the $\mu$ dependence on $C(t)$ factorizes as well.
We now turn to the finite temperature case. The proof of factorization of the $\mu$ dependence at finite temperature, presented in Appendix A, follows along the same lines as for $\beta=0$. A finite-temperature steady state $\rho=\exp(-\beta H)/Z$, with $Z={\rm Tr}\exp(-\beta H)$ can be prepared by a judicious choice of jump operators with temperature-dependent couplings [64] (see Appendix A). In this case, we define the quantum uncertainty as
$C(t,\beta,\mu)$ $=$ $\sum_{i\neq j}\frac{\langle\{\chi_{i}(t)\rho^{1/4},\chi_{j}\rho^{1/4}\}^{2}\rangle}{N(N-1)}$ (8)
$=$ $e^{-2\mu t}(F(t,\beta)+I(t,\beta))$
$\equiv$ $e^{-2\mu t}C(t,\beta)$
where $F(t,\beta)$ and $I(t,\beta)$ are defined with the corresponding $\rho^{1/4}$ insertions [18] with respect to the $\beta=0$ analogues. As a consequence, it is only necessary to compute analytically $C(t,\beta)$ for a Hermitian $q=2$ SYK model at finite temperature.
In order to proceed, we choose a basis in which $H$ is diagonal,
$H=i\sum_{k=1}^{N/2}\lambda_{k}\widetilde{\chi}_{2k-1}\widetilde{\chi}_{2k},$ (9)
where $\widetilde{\chi}_{i}=S_{ik}\chi_{k}$, with $S_{ik}$ an orthogonal matrix, and $\lambda_{k}$ are the eigenvalues of the antisymmetric real couplings $J_{ij}$. In this new basis, the OTOC is given by
@@
$\times\left\langle\int dSS^{T}_{1k}S^{T}_{2m}S^{T}_{1p}S^{T}_{2q}\right\rangle.$
It is now necessary to carry out the averages over the orthogonal matrices $S_{ij}$. This is possible by employing the relation [65]
$\langle S_{1k}S_{2m}S_{1p}S_{2q}\rangle=\frac{N+1}{\mathcal{N}}\delta_{kp}\delta_{mq}-\frac{1}{\mathcal{N}}(\delta_{km}\delta_{qp}+\delta_{kq}\delta_{mp}),$
where ${\cal N}=N(N-1)(N+2)$, which results in three different contributions to $F(t,\beta)$. The first two are equal because of the reflection symmetry of the spectrum. The OTOC can be further simplified using $\widetilde{\chi}_{k}^{2}=1/2$ after commuting the four Majorana operators through the evolution operator and the density matrix. The resulting sums are expressed in terms of $\lambda_{k}$ by using the properties of the trace over the many-body states. The final finite-$N$ result for $C(t,\beta)$ Eq. (8) is
@@
III Analytic expressions to orders $1/N$ and $1/N^{2}$
The above expressions simplify by keeping only the leading $1/N$ correction, replacing the sums with integrals and noticing that the spectral density of the eigenvalues $\lambda_{k}$ is $\rho(\lambda)=(1/\pi)\sqrt{1-\lambda^{2}/4}$ [66]. For simplicity, we henceforth restrict ourselves to $\beta=0$. Considering only the leading $1/N$ correction, $C(t,0)\equiv C(t)$ Eq. (8), with $F(t,0)$ and $I(t,0)$ given by Eqs. (12), (13), simplifies to
@@ III Analytic expressions to orders $1/N$ and $1/N^{2}$
where $J_{1}(t)$ is a Bessel function of first kind and time is measured in units of $J=1$. The upper plot in Fig. 1 demonstrates the convergence of the numerical results of $C(t)$ obtained from Eqs. (12) and (13) to the analytic prediction Eq. (15). Analogous analytic results for finite $\beta$ are rather cumbersome. In Appendix A, we have checked that finite temperature effects do not change qualitatively the dynamics, at least for sufficiently long times. A consequence of Eq. (15) is that, for $t\ll 1$, $C(t)$ grows quadratically as $t^{2}$. For $t\gg 1$, the system approaches a steady state in a power-law fashion with superimposed oscillations. The effect of the bath is an overall exponential decay of $C(t)$ that, as expected, suppresses quantum scrambling. For a sufficiently weak coupling, $\mu\ll 1$, the system approaches the steady state in two stages: first, the mentioned power-law, and only for longer times $\sim 1/\mu$ the exponential suppression of $C(t)$.
Figure 1: Time evolution of the quantum uncertainty $C(t)$ with $t$ in units of $J=1$ for $\mu=\beta=0$. Top: The blue ($N=20$), red ($N=200$), and green ($N=2000$) curves correspond to $C(t)$ given in Eq. (12) and (13) and $C_{N}=(N+1)/[(N-1)(N+2)]$. They converge to the $1/N$ prediction Eq. (15), with $C_{N}=1/N$, only for sufficiently large NN. In both cases, $C_{N}$ is given by the time independent part of $C(t)$. Bottom: Comparison between the analytical exact $C(t)$, Eqs. (8), (12) and (13), in terms of the eigenvalues of $J_{ij}$ (purple circles) with the simple analytic $C(t)$, Eq. (III), including up to $1/N^{2}$ corrections for $N=200$ on a linear scale. The black curve stands for the $1/N$ result, Eq. (15), while the green curve stands for the $1/N+1/N^{2}$ result, Eqs. (III) and (19). The agreement between of analytical exact and the $1/N+1/N^{2}$ results is excellent for all timescales. Deviations from the $1/N$ expression occur at $t\sim\sqrt{N}$. The Heisenberg time, signaling saturation, is at $t\sim 2N$.
@@
where we have reinstated $\cal N$ in order to account for additional $1/N^{2}$ effects, $\theta$ is the Heaviside step function, and
$N^{2}K_{c}(t)=\langle|\sum_{k=1}^{N}\exp(it\lambda_{k})|^{2}\rangle-|\langle\sum_{k=1}^{N}\exp(it\lambda_{k})\rangle|^{2}$ (17)
is the connected spectral form factor (SFF).
@@
For times $t\gtrsim\sqrt{N}$, $K_{c}(t)\sim 2t/\pi N$ and then gradually saturates to $1/N$ as the Heisenberg time $t\sim 2N$ is approached. A comment is in order. This SFF is computed from the eigenvalues of the couplings $J_{ij}$, so in line with previous results [69, 70], it can be interpreted as the one obtained from the single-particle quantum chaotic motion. Therefore, we do not observe an exponential ramp as the one claimed in Refs. [71, 72, 73]. We stress that since the spectral density is not constant, the time dependence of Eq. (19) has clear deviations from linear behavior before the Heisenberg time, which translates into similar deviations in $C(t)$. Interestingly, in the range of times of interest, $2N>t>\sqrt{N}$, a simple inspection of Eq. (III) reveals that $C(t)$ is controlled by $K_{c}(t)$ despite it being of the order $1/N^{2}$ and therefore subleading in the $1/N$ expansion. The reason for that is that the time dependence of the leading $1/N$ terms approaches zero in a power-law fashion so that for $t>\sqrt{N}$ they become smaller than $K_{c}(t)$ which increases linearly in time. Crucially, the sign of the SFF term is negative, which leads to a local maximum in the OTOC not related to the small oscillations. Therefore, the OTOC decreases with time for $t>\sqrt{N}$, which corresponds to a decrease of quantum scrambling due to the discreteness of the spectrum captured by the SFF.
@@
The results depicted in the bottom panel of Fig. 1 fully confirm this picture. The analytical expression for $C(t)$ Eq. (III), which includes $1/N$ and $1/N^{2}$ corrections, agrees at all times scales with the exact $C(t)$ Eqs. (8), (12), (13), at $\beta=0$, in terms of the eigenvalues of $J_{ij}$, which must be computed numerically. Substantial deviations are observed for long times if only the leading $1/N$ correction Eq. (15) is taken into account. The effect of the bath is to suppress scrambling exponentially.
@@
We have thus found that, for $t\gtrsim\sqrt{N}$, $C(t)$ is controlled by the SFF of the random couplings $J_{ij}$, which has the expected linear behavior observed in quantum chaotic systems for times $t>\sqrt{N}$ but not too close to the Heisenberg time. This does not mean that the $q=2$ SYK is many-body quantum chaotic. While single-particle observables may show quantum chaos, the many-body dynamics is still integrable because the Majoranas are not interacting. For instance, the level statistics is Poisson. Moreover, the timescales involved are polynomial in $N$ which is typical of integrable systems. The results of Appendix A show that finite $\beta$ effects do not change qualitatively this picture. Before we embark on the comparison with numerical results, we comment on the differences and similarities between the $C(t)$ of a quantum chaotic system and that of the integrable SYK model. Starting with the similarities, the perturbative quadratic growth and the power-law decay have also been observed in quantum chaotic systems [69, 29, 70, 74]. Moreover, the observed non-monotonicity is also expected in quantum chaotic systems because it is a general consequence of the discreteness of the spectrum. Finally, at least for short times [50], the effect of the bath induces a similar decay in both cases. However, there are also differences, for instance, the OTOC for the integrable SYK model experiences neither the exponential growth around the Ehrenfest time nor the exponential decay characterized by the Ruelle-Pollicott resonances [3, 11] as the system approaches thermalization, which are features in the quantum chaotic case. Finally, the saturation time is $\sim\exp(N/2)$ for quantum chaotic many-body systems while it is linear $\sim N$ for the integrable SYK model.
@@
Figure 2: Top: $C(t)$ with $t$ in units of $J=1$, for $\mu=\beta=0$. $C(t)$ Eq. (8) is computed using exact diagonalization for $N=24$ (red) with $\sim 10^{6}$ disorder realizations. The exact diagonalization result shows almost no difference with respect to the exact analytic result (black), Eqs. (12) and (13), and and also an excellent agreement with the $1/N+1/N^{2}$ simple expression (blue) given in Eq. (III). Substantial deviations are observed if only the leading $1/N$ correction (green) Eq. (15) is kept. We are neglecting terms $1/N^{2}$ in $NC(t)$. Since $N=24$ and, at least to lower order, prefactors are of the order one, differences between ED and the analytic $1/N+1/N^{2}$ expression (blue curve) should be $\sim 1/24^{2}\approx 0.0017$, is fully consistent with the observed deviation. Bottom: $C(t)$ for $\mu=0.15$, $N=18$ and $\beta=0$. We combine exact diagonalization with the quantum trajectory method [75, 76]. For each point in time, we use 200200 quantum trajectories, a time step of $dt=0.01$, and at least $5\times 10^{5}$ disorder realizations. The exponential approach to the steady state for $\mu\neq 0$ is correctly reproduced by the quantum trajectory method and we find good agreement with the $1/N$ semiclassical result.
IV Exact numerical results